OpenStudy (anonymous):

How do you factor 3x^4-3? Can you please help me? I need to answer this. Thankyou.

hartnn (hartnn):

can you first factor out 3 from 3x^4-3 ?

OpenStudy (mathstudent55):

First always look for a common factor. Then look at number of terms: If 2 terms, try difference of squares or sum or difference of cubes.

OpenStudy (anonymous):

hartnn, thankyou! I'll try that!

hartnn (hartnn):

good! if you get stuck, we'll be here to help :)

OpenStudy (anonymous):

Thankyou!! I always need help with my Math homework XD I'm not that good but I am willing to learn.

hartnn (hartnn):

oh, and bdw, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

Thankyou! That's very nice of you :">

OpenStudy (anonymous):

Is the answer 3(x^2+1)(x^2-1) ?

hartnn (hartnn):

you are on right path! but notice that you can further factor x^2-1 :)

OpenStudy (mathstudent55):

So far, so good, but after you factor, you need to check each factor to see if you can factor some more.

OpenStudy (anonymous):

So I have to factor it again? I'm kinda stuck XDD

hartnn (hartnn):

same thing what you have done, \(a^2-b^2= (a+b)(a-b) \\ (x^2-1^2) =... ?\)

OpenStudy (mathstudent55):

Look at each factor, and try this again: If 2 terms, try difference of squares or sum or difference of cubes.

OpenStudy (anonymous):

Thankyou both! :)) I'll try that right now! Thanks. You're both great help!

OpenStudy (anonymous):

For x^2-1 is it (x+1)(x-1) ?

hartnn (hartnn):

correct :)

OpenStudy (anonymous):

Thankyou. I still have to figure out x^2 + 1. I'm stuck again XDD. haha.

hartnn (hartnn):

now are you aware of imaginary numbers ? [if not, keep x^2+1 as it is] or does this belong to imaginary/complex numbers ?

OpenStudy (anonymous):

I'm not really aware of it... Sorry

hartnn (hartnn):

x^2+1 does not have real factors, here, lets assume real factorization so, your final answer will be \(\large 3 (x+1)(x-1)(x^2+1)\) thats it!

OpenStudy (anonymous):

Thankyou so much

hartnn (hartnn):

you are most welcome ^_^

OpenStudy (mathstudent55):

\[3x^4 - 3\] \[=3(x^4 - 1)\] \[=3(x^2 + 1)(x^2 -1)\] \[=3(x^2 + 1)(x + 1)(x - 1)\]

hartnn (hartnn):

if and when you learn imaginary numbers, x^2+1 will be factorized as \((x+i)(x-i)\) till then , you can ignore this :)

OpenStudy (anonymous):

Thankyou both! I'm moving on to the next problem. Thankyousomuch for helping me with this. :))

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

x^6 + 26x^3 - 27 is it equal to : (x^3 + 27)(x^3 - 1) ?

OpenStudy (mathstudent55):

Yes, but you're not done with factoring. Look at both binomials.

OpenStudy (mathstudent55):

If 2 terms, try difference of squares or sum or difference of cubes.

OpenStudy (mathstudent55):

@MoeshaKiel Did you finish this problem?

OpenStudy (anonymous):

Yes. I did. Thanks to the both of you :)) you were awesome help!

OpenStudy (mathstudent55):

ok, wlcm